# Large Language Model-Brained GUI Agents: A Survey

### v9

@@ 11.5 Customization and Personalization
For GUI agents to be truly effective, they need to understand individual user preferences and adapt their behavior accordingly [45, 503]. Users have unique habits, workflows, and preferences, and a one-size-fits-all approach may not provide an optimal user experience [504]. For example, an agent that assists with document editing may need to learn a user’s preferred writing style, formatting choices, and commonly used phrases. Without personalization, the agent may suggest edits that the user finds unhelpful or inconsistent with their style. Customization requires the agent to learn and adapt to user-specific preferences [505], which can be challenging due to variability among users. Balancing personalization with privacy concerns is also critical, as collecting and utilizing personal data must be handled responsibly.
@@ 11.6 Ethical and Regulatory Challenges
The deployment of LLM-powered GUI agents raises ethical questions regarding accountability, transparency, fairness, and user consent [454, 512, 513, 514]. There is a risk of biased behavior inherited from training data, leading to unfair or discriminatory actions [515, 516]. For example, an agent used in hiring processes might inadvertently exhibit bias by filtering out resumes based on gender or ethnicity if such biases are present in the training data. Additionally, regulatory compliance varies across industries and regions, complicating deployment.
@@ 11.7 Scalability and Generalization
Many existing GUI agents are tailored to specific applications or environments, which limits their scalability and generalizability. The diversity of software interfaces, with each application featuring unique designs, layouts, and interaction patterns, poses a significant challenge for developing agents capable of operating seamlessly across multiple platforms, even for common pop-up windows [521]. For instance, an agent developed to automate tasks in a specific version of a word processor may fail when the application updates its interface or when used with a different word processor that has a different UI layout. This issue becomes even more pronounced when the agent encounters applications or environments that it is unfamiliar with or has not encountered during training. Even if these new environments share some similarities with previously seen GUIs, the agent may still make mistakes and require exploration to fully understand their functionality. The lack of generalization limits the agent’s applicability and necessitates continuous updates or retraining, which can be resource-intensive [522, 523].

### v11

@@ 11.5 Customization and Personalization
Effective GUI agents must go beyond generic task completion and provide experiences that are personalized to individual users, adapting to their unique workflows, preferences, and behavioral patterns [45, 595]. Unlike general LLM applications that operate in isolated prompts or conversations, GUI agents work across software environments where user interaction styles can vary significantly. A one-size-fits-all agent may fail to align with how a particular user edits documents, navigates interfaces, or organizes tasks—resulting in friction, inefficiency, or user frustration [596].
For instance, a GUI agent assisting with document editing must learn the user’s preferred tone, formatting conventions, and vocabulary. Without this contextual understanding, the agent may offer irrelevant suggestions or enforce formatting inconsistent with the user’s intent. Personalization in GUI agents thus requires longitudinal learning, where the agent continually adapts based on prior interactions, fine-tunes its behavior to match user expectations, and preserves consistency across sessions [597].
However, this introduces new challenges. The high variability in user preferences—especially in free-form GUI environments—makes it difficult to define universal personalization strategies. Moreover, collecting and leveraging user-specific data must be done responsibly, raising critical concerns around privacy, data retention, and on-device learning. Striking a balance between effective customization and user trust is particularly important for GUI agents, which often operate over sensitive documents, personal applications, or system-level interfaces.
@@ 11.6 Ethical and Regulatory Challenges
LLM-powered GUI agents raise distinct ethical and regulatory concerns due to their ability to perform real-world actions across software interfaces. Unlike traditional LLMs, these agents can autonomously trigger operations, manipulate data, and interact with sensitive applications—amplifying risks around accountability, fairness, and user consent [539, 604, 605, 606, 607].
A key concern is bias inherited from training data, which can lead to unfair behavior in sensitive workflows. For example, a GUI agent assisting in hiring may unknowingly exhibit gender or racial bias [608, 609]. These risks are harder to audit at the GUI level due to limited traceability across multi-application actions. Regulatory compliance adds further complexity. GUI agents often operate across domains with strict data protection laws, but lack standardized mechanisms for logging actions or securing user consent. This makes it challenging to meet legal and ethical standards, especially when agents act in opaque or background contexts. Addressing these issues requires tailored solutions for GUI agents, including permission controls, runtime confirmations, and transparent activity logs—ensuring safe, fair, and compliant deployment across diverse environments.
@@ 11.7 Scalability and Generalization
GUI agents often struggle to scale beyond specific applications or environments, limiting their generalization. Each software interface features unique layouts, styles, and interaction patterns—even common UI elements like pop-up windows can vary widely [614]. These variations make it difficult to design agents that operate robustly across platforms without retraining or fine-tuning.
A further challenge is the dynamic nature of real-world GUIs. Frequent changes due to software updates, A/B testing, or interface redesigns—such as repositioned buttons or modified widget hierarchies—can easily break previously functional agents. For example, an agent trained on one version of a word processor may fail when the layout changes, or when deployed on a different program with similar functionality but a different interface structure. Even when GUIs share visual similarities, agents often fail to generalize without additional exploration or adaptation [615]. This lack of robustness restricts deployment in practical settings and increases the cost of maintenance, requiring frequent updates or retraining to stay aligned with evolving environments [616, 617, 618]. Overcoming this challenge remains critical for developing truly scalable and adaptable GUI agents.